# Benchmarks and Challenges in Pose Estimation for Egocentric Hand Interactions with Objects

### v1

@@ Abstract
Equal contribution, 1ETH Zürich, 2Max Planck Institute for Intelligent Systems, Tübingen, Germany, 3The University of Tokyo, 4Communication University of China, 5Jiiov, 6Bytedance, 7Fujitsu Research & Development Center Co., Ltd., 8Beihang University, 9RWTH Aachen University, 10UNIST, 11UIUC, 12Meta, 13University of Birmingham, 14NUS.
@@ 1 Introduction
Figure 1: Tasks in HANDS23 based on AssemblyHands and ARCTIC. In AssemblyHands, from its multi-view headset (a), the goal is to estimate 3D hand poses from images (b); In ARCTIC, given an image, the goal is to estimate the poses of two hands and articulated object surface models (c).
@@
We interact with the world with our hands and see it through our eyes: we wake up and grab our phone to check the time; we use tools when assembling parts of a car; we open the microwave door to heat up food, to name a few. An egocentric 3D understanding of our hand interactions with objects will fundamentally impact areas such as robotics grasping [74, 10], augmented and virtual reality [22], action recognition [18, 66, 6] and motion generation [74, 73].
@@
However, it is non-trivial to accurately reconstruct 3D hands and/or objects due to its high degree of freedoms [67, 59], ambiguous texture [15], and heavy occlusions. These challenges are intensified in an egocentric view [47], particularly with object interactions, due to significant camera distortion, rapid and varied changes in viewpoint caused by head movements, and hand-object occlusion. To better understand these challenges, we introduce a public challenge in conjunction with ICCV 2023 (i.e., HANDS23) based on recent egocentric hand datasets, AssemblyHands [46] and ARCTIC [16] (see Figure 1). These two datasets are large-scale, multi-view, and provide monochrome or RGB egocentric videos of the hands dexterously manipulating objects. Accordingly, we host two tasks: 1) egocentric 3D hand pose estimation from a single-view image based on AssemblyHands, and 2) consistent motion reconstruction based on ARCTIC.
@@
We introduce new methods from HANDS23 as well as recent dataset leaderboard baselines that substantially outperform initial baselines for both tasks, setting new benchmarks for subsequent comparisons on the datasets. With these benchmarks, we thoroughly analyze factors such as viewpoint, action types, hand position, model size, and object variations, to determine their impact on 3D hand(-object) reconstruction.
@@
Our findings show the success of addressing the distortion of egocentric cameras with explicit perspective cropping or implicit learning for the distortion bias. In addition, recent high-capacity vision transformers are capable of learning complex hand-object interactions. Adaptive fusion techniques for multi-view predictions further boost performance. We also analyze the remaining challenges that are still difficult to handle with the recent methods, e.g., fast hand motion, object reconstruction from narrow and moving views, and intricate interactions and close contact between two hands and objects.
@@
To summarize, we contribute state-of-the-art baselines and gather the submitted methods for AssemblyHands and ARCTIC to foster future research on egocentric hand-object interactions. Furthermore, we thoroughly analyze the two benchmarks to provide insights for future directions in the egocentric hand pose estimation task and the consistent motion reconstruction task.
@@ 2 Related work
3D hand pose estimation: Reconstructing 3D hand poses has a long history [13, 45] ever since the important work led by Rehg and Kanade [52]. A large body of research in this area focuses on single-hand reconstruction [29, 43, 60, 3, 19, 58, 59, 80, 2, 25, 79, 15, 57, 75, 65, 8, 17, 5, 35, 12]. For example, while a popular OpenPose library features 2D hand keypoints  [57], Zimmermann et al. [80] initially extend to estimate 3D hand poses using deep convolutional networks. Ever since the release of InterHand2.6M [42] dataset, the community has increased focus on strongly-interacting two hands reconstruction [31, 40, 32, 34, 39, 42, 41, 33, 46, 21]. For example, Li et al. [33] and Moon et al. [41] use relighting techniques to augment InterHand2.6M with more natural lighting and diverse backgrounds.
@@
Recently, there are methods that do not assume object templates for hand-object reconstruction [25, 69, 9, 14, 70, 28, 62]. For example, Fan et al. [14] introduced the first category-agnostic method that reconstructs an articulated hand and object jointly from a video. However, here we focus on template-based approaches because challenges and insights in the template-based approaches such as hand-object occlusion, camera distortion are still transferable to more challenging template-free reconstruction settings.
Public reports for the previous challenges (HANDS17 [71] and HANDS19 [1]) have acted to distill the insights from individual review papers and practical techniques into comprehensive summaries, thereby enriching the community’s knowledge base. These past challenges have established benchmarks, each of which includes depth-based hand pose estimation from egocentric views. Instead of depth sensors, the HANDS23 benchmarks are based on affordable and widely applicable image sensors, i.e., RGB and monochrome images. Additionally, this paper advances the analysis further with unique insights, such as multi-view egocentric cameras, object reconstruction in contact, and modeling with recent transformers beyond conventional CNNs.
@@ 3.1 Workshop challenges
3D hand pose estimation in AssemblyHands: As illustrated in Figure 1, this task focuses on egocentric 3D hand pose estimation from a single-view image based on AssemblyHands. The dataset provides multi-view captured videos of hand-object interaction during assembling and disassembling of toy-vehicles. In particular, it provides both allocentric and egocentric recordings, and auxiliary cues like action, object, or context information for hand pose estimation. We refer readers to [46] for more dataset details. The training, validation, and testing sets contain 383K, 32K, and 62K monochrome images captured from egocentric cameras, respectively. During training, 3D hand keypoint coordinates, hand bounding boxes and camera intrinsic and extrinsic matrix for four egocentric cameras attached to the headset are provided. During testing, all annotations except for the 3D keypoints are provided. Unlike the original work [46], given the availability of multi-view egocentric images, this task lets participants develop multi-view fusion based on the corresponding multi-view images.
@@ 3.2 Evaluation criteria
AssemblyHands evaluation: We use mean per joint position error (MPJPE) as an evaluation metric in millimeters, comparing the model predictions against the ground-truth in world coordinates. We provide the intrinsic and extrinsic camera parameters of the egocentric cameras to construct submission results defined in the world coordinates. Assuming that the human hand has a total ${N}_{J}$ joints, we denote wrist-relative coordinates of the prediction and ground-truth as $\hat{J}\in\mathbb{R}^{{N}_{J}\times 3}$ and $J\in\mathbb{R}^{{N}_{J}\times 3}$, respectively. Given a joint visibility indicator $\gamma_{i}$ per joint $J_{i}$, we compute the Euclidean distance between predicted and ground-truth joints as $\frac{1}{\sum_{i=1}^{N_{J}}\gamma_{i}}\sum_{i=1}^{N_{J}}\gamma_{i}\left\|\hat{J}_{i}-J_{i}\right\|_{2}$.
@@
For completeness, we report all metrics introduced in ARCTIC. In particular, the task requires the reconstructed meshes to have accurate hand-object contact (CDev), and smooth motion (ACC). Additionally, during articulation or when carrying an object, it’s crucial that vertices of the hand and object in constant contact maintain synchronized movement (MDev). Moreover, we assess hand and object poses, alongside their relative movements, using metrics like MPJPE, AAE, Success Rate, and MRRPE. For detailed information, see [16].
@@ 4 Methods
In this section, we present the methods in the two challenges as well as other competing methods on the leaderboards. In total, there are four methods outperforming the baseline in both AssemblyHands and ARCTIC, which will be the focus in this paper.
@@ 4.1 AssemblyHandsmethods
Participants develop methods that learn the mapping from egocentric images to 3D keypoints. The methods are categorized into: heatmap-based and regression-based approaches. Given the presence of complex hand-object interactions in the egocentric scenes, high-capacity transformer models and attention mechanisms addressing occluded regions have been proposed as the backbone networks. Table 1 summarizes the methods in terms of learning methods, preprocessing, multi-view fusion, and post-processing.
@@
JHands: In contrast to the baseline, this method employs a regression-based approach with simple MLP heads for regressing 2D keypoints, root-relative 3D keypoints, and the global root depth. The regression training is empowered by a recent fast and strong vision transformer architecture, Hiera [55], pre-trained with masked auto-encoder [27]. A multi-level feature fusion operation that concatenates the features of different layers is adopted for better feature extraction at different scales. The method additionally uses other publicly available datasets for training, namely FreiHAND [80], DexYCB [5], and CompHand [7].
@@
Table 1: Method summary in AssemblyHands. We summarize submitted methods in terms of learning methods, architecture, preprocessing, and multi-view fusion techniques. The tuple (views, phase) indicates the number of views used in either train or test time.
@@
JHands Regression Hiera [55]
@@
FRDC: This method adopts a hybrid approach by combining regression with heatmap for training. HandOccNet [49] is modified to regress 3D keypoint coordinates and integrated with an additional branch of 2D heatmap regression. HandOccNet enriches feature extraction for occluded regions with spatial attention mechanisms, making it robust under hand-object occlusions. The method further utilizes a stronger ConvNeXt [38] backbone and feature fusion from the 2D keypoint regressor.
@@
While Base uses a simple average of predicted keypoints from all four camera views in the test time, PICO-AI proposes multi-view feature fusion during training using Feature Transform Layers (FTL) [53]. This FTL training requires fusing two out of four views; thus, the method chooses the most suitable views for every frame. If multiple candidates exist, it computes the Intersection over Union (IoU) between hand boxes generated from per-view prediction and 2D keypoints projected from the previous 3D predictions. Then, the two views with the highest IoUs are selected for their superior prediction reliability compared to the rest.
The methods JHands, FRDC, and Phi-AI apply adaptive fusion in predicted keypoints during testing. The method JHands computes the MPJPE with each other view and selects two results of views that have the lowest MPJPE, excluding noisy predictions in the average. If the MPJPE is lower than a threshold, the mean of the two results is calculated as the final result. Otherwise, the result that has a lower PA-MPJPE with the predictions in the previous frame is chosen. The methods FRDC and Phi-AI use a weighted average for each view prediction, assigning weights based on each view’s validation performance.
Postprocessing: Several postprocessing techniques, including test-time augmentation, smoothing, and model ensemble, have been used to enhance inference outcomes. In particular, the method JHands applies an offline smooth (Savitzky-Golay) filter on each video sequence.
@@ 4.2 ARCTICmethods
Table 2: Method summary in ARCTIC. We summarize baselines on ARCTIC benchmark in terms of input dimensions, image backbones, learning rate scheduling, training epochs, batch size and the cropping used for input. Method trains 50 epochs for decoder and 36 for backbone. +Learning rate is 1e-7 to 1e-4 with linear warmup for first 5% step, and 1e-4 to 1e-7 with cosine decay for rest.
@@
ArcticNet-SF [16]: Introduced in ARCTIC, it is a single-frame baseline. It first extract a image feature vector from the input image, then it regresses hand and object parameters with simple MLPs. The hand and object meshes can then be extracted via $\mathcal{H}(\cdot)$ and $\mathcal{O}(\cdot)$. For more details, see [16].
JointTransformer [30]: JointTransformer enhances ArcticNet-SF by integrating a transformer decoder in place of the MLP regressors for hand and object parameter estimation. The decoder employs learned queries for the angle of each joint, as well as the shape and translation of each hand and the translation, rotation, and articulation of the object. It alternates between self-attention between queries and cross-attention of queries to the elements of the backbone feature map, followed by linear layers that regress the final parameters. Specifically, there are separate linear layers dedicated to regressing joint angles, hand shape, hand translation, object translation, object orientation, and object articulation. The best model uses a ViT-G [72] backbone with frozen DINOv2 weights [48].
@@
UVHand: Since ArcticNet-SF only leverages a global feature vector to estimate hand and object parameters, the image features lack local context. To address this, UVHand leverages Swin-L transformer [37] to extract image features. They then further leverage Deformable DETR [78] to encode the feature maps in multiple scales. The encoded feature maps are then being aggregated via self and cross attention and finally being used to regress hand and object parameters.
DIGIT [15]: The method was introduced to estimate strongly interacting hands in [15]. Since the ArcticNet-SF have sensitive prediction when hands interacting with objects, DIGIT was extended to the ARCTIC setting. Given an image, it first estimate hand part segmentation masks and object mask. The mask predictions are fused with the image features to perform parameter estimation.
Implementation details: As showned in Table 2, all methods use the default cropping method as in ARCTIC to crop around the object while AmbiguousHands performs three crops (around two hands and the object). DIGIT uses the HRNet-W32 backbone [61] and train with a batch size of 64 with the same learning rate for all iterations. UVHand takes as input a $384\times 384$ image cropped around the object, encodes it with the Swin-L transformer [37] backbone. Their method was trained with a batch size of 48 with a learning rate of 2e-4 for the backbone and 1e-7 for other weights. Due to computational cost, they train 50 epochs for the decoder and 36 for the backbone. JointTransformer uses ViT-G [72] backbone with frozen DINOv2 [48] weights to train with a batch size of 64. It performs a linear warmup from 1e-7 to 1e-4 in the first 5% steps and uses cosine decay from 1e-4 to 1e-7 for the rest of the steps.

### v2

@@ Abstract
Equal contribution, 1ETH Zürich, 2Max Planck Institute for Intelligent Systems, Tübingen, Germany, 3The University of Tokyo, 4Communication University of China, 5Jiiov, 6Bytedance, 7Fujitsu Research & Development Center Co., Ltd., 8Beihang University, 9RWTH Aachen University, 10UNIST, 11UIUC, 12Meta, 13University of Birmingham, 14National University of Singapore.
@@ 1 Introduction
Figure 1: Tasks in HANDS23. In AssemblyHands, from its multi-view headset (a), we estimate 3D hand poses from images (b); In ARCTIC, we estimate the poses of two hands and articulated objects from an image (c).
@@
We interact with the world with our hands and see it through our eyes: we wake up and grab our phone to check the time; we use tools when assembling parts of a car; we open the microwave door to heat food, to name a few. An egocentric 3D understanding of our hand interactions with objects will fundamentally impact areas such as robotics grasping [80, 10], augmented and virtual reality [24], action recognition [18, 72, 6] and motion generation [80, 79].
@@
However, it is non-trivial to accurately reconstruct 3D hands and or objects due to its high degree of freedoms [73, 65], ambiguous texture [15], and heavy occlusions. These challenges are intensified in an egocentric view [50, 51], particularly with object interactions, due to significant camera distortion, rapid and varied changes in viewpoint caused by head movements and hand-object occlusion. To better understand these challenges, we introduce a public challenge in conjunction with ICCV 2023 (i.e., HANDS23) based on recent egocentric hand datasets, AssemblyHands [49] and ARCTIC [16] (see Figure 1). These two datasets are large-scale, multi-view, and provide monochrome or RGB egocentric videos of the hands dexterously manipulating objects. Accordingly, we host two tasks: 1) egocentric 3D hand pose estimation from a single-view image based on AssemblyHands, and 2) consistent motion reconstruction based on ARCTIC.
@@
We introduce new HANDS23 methods and recent dataset leaderboard baselines that substantially outperform initial baselines for both tasks, setting new benchmarks for subsequent comparisons on the datasets. The two datasets are significantly larger and include a wider variety of bimanual manipulations compared to earlier datasets [22, 5], enabling a more authentic assessment of real-world interactions. With these benchmarks, we thoroughly analyze factors such as viewpoint, action types, hand position, model size, and object variations to determine their impact on 3D hand(-object) reconstruction.
@@
Our findings show the success of addressing the distortion of egocentric cameras with explicit perspective cropping or implicit learning for the distortion bias. In addition, recent high-capacity vision transformers can learn complex hand-object interactions. Adaptive fusion techniques for multi-view predictions further boost performance. We also analyze the remaining challenges that are still difficult to handle with the recent methods, e.g., fast hand motion, object reconstruction from narrow and moving views, and intricate interactions and close contact between two hands and objects.
@@
To summarize, we contribute state-of-the-art baselines and gather the submitted methods for AssemblyHands and ARCTIC to foster future research on egocentric hand-object interactions. Furthermore, we thoroughly analyze the two benchmarks to provide insights for future directions in egocentric hand pose estimation and consistent motion reconstruction.
@@ 2 Related work
3D hand pose estimation: Reconstructing 3D hand poses has a long history [13, 48] ever since the important work led by Rehg and Kanade [57]. A large body of research in this area focuses on single-hand reconstruction [31, 46, 66, 3, 19, 64, 65, 88, 2, 27, 87, 15, 63, 81, 71, 8, 17, 5, 37, 12]. For example, while a popular OpenPose library features 2D hand keypoints  [63], Zimmermann et al. [88] initially extend to estimate 3D hand poses using deep convolutional networks. Ever since the release of InterHand2.6M [45] dataset, the community has increased focus on strongly interacting two hands reconstruction [33, 43, 34, 36, 42, 45, 44, 35, 49, 21]. For example, Li et al. [35] and Moon et al. [44] use relighting techniques to augment InterHand2.6M with more natural lighting and diverse backgrounds.
@@
More recent methods do not assume object templates for hand-object reconstruction [27, 75, 9, 14, 76, 30, 68]. For example, Fan et al. [14] introduced the first category-agnostic method that reconstructs an articulated hand and object jointly from a video. Nonetheless,our challenges and insights on hand-object occlusions and camera distortions are still applicable to these more challenging template-free reconstruction settings.
Public reports for the previous challenges (HANDS17 [77] and HANDS19 [1]) distilled the insights from individual review papers and practical techniques into comprehensive summaries to enrich the community’s knowledge base. These past challenges use benchmarks which include depth-based hand pose estimation from egocentric views. Instead of depth sensors, the HANDS23 benchmarks are based on affordable and widely applicable image sensors, i.e., RGB and monochrome images. This paper further advances the analysis with unique insights, such as multi-view egocentric cameras, object reconstruction in contact, and modeling with recent transformers beyond conventional CNNs.
@@ 3.1 Workshop challenges
3D hand pose estimation in AssemblyHands: As illustrated in Figure 1, this task focuses on egocentric 3D hand pose estimation from a single-view image based on AssemblyHands. The dataset provides multi-view captured videos of hand-object interaction while assembling and disassembling toy vehicles. In particular, it provides allocentric (fixed-view) and egocentric recordings and auxiliary cues like action, object, or context information for hand pose estimation. We refer readers to [49] for more dataset details. The training, validation, and testing sets contain 383K, 32K, and 62K monochrome images, respectively, all captured from egocentric cameras. During training, 3D hand keypoint coordinates, hand bounding boxes and camera intrinsic and extrinsic matrices for the four egocentric cameras attached to the headset are provided. The same information is provided during testing, minus the 3D keypoints. Unlike [49], given the availability of multi-view egocentric images, this task lets participants develop multi-view fusion based on the corresponding multi-view images.
@@ 3.2 Evaluation criteria
AssemblyHands evaluation: We use MPJPE as an evaluation metric in millimeters, comparing the model predictions against the ground-truth in world coordinates. We provide the intrinsic and extrinsic of the egocentric cameras to construct submission results defined in the world coordinates. Assuming that the human hand has a total ${N}_{J}$ joints, we denote wrist-relative coordinates of the prediction and ground-truth as $\hat{J}\in\mathbb{R}^{{N}_{J}\times 3}$ and $J\in\mathbb{R}^{{N}_{J}\times 3}$, respectively. Given a joint visibility indicator $\gamma_{i}$ per joint $J_{i}$, we compute the Euclidean distance between predicted and ground-truth joints as $\frac{1}{\sum_{i=1}^{N_{J}}\gamma_{i}}\sum_{i=1}^{N_{J}}\gamma_{i}\left\|\hat{J}_{i}-J_{i}\right\|_{2}$. The visibility indicator offers per-joint binary labels, representing whether the annotated keypoints are visible from the given egocentric view.
@@
For completeness, we report all metrics introduced in ARCTIC. In particular, the task requires the reconstructed meshes to have accurate hand-object contact (CDev) and smooth motion (ACC). Additionally, during articulation or when carrying an object, it is crucial that the vertices of the hand and object which are in contact maintain synchronized movement (MDev). Moreover, we assess hand and object poses, alongside their relative movements, using metrics like MPJPE, AAE, Success Rate, and MRRPE. For detailed information, see [16].
@@ 4 Methods
This section presents the methods in the two challenges and other competing methods on the leaderboards. Four methods outperform the baseline in both AssemblyHands and ARCTIC.
@@ 4.1 AssemblyHandsmethods
Participants develop methods that learn the mapping from egocentric images to 3D keypoints. The methods are categorized into: heatmap-based and regression-based approaches. Given the presence of complex hand-object interactions in the egocentric scenes, high-capacity transformer models and attention mechanisms addressing occluded regions have been proposed as the backbone networks. Table 1 summarizes the methods based on the learning, preprocessing, multi-view fusion, and post-processing approaches.
@@
JHands [85]: This method employs a regression-based approach with simple MLP heads for regressing 2D keypoints, root-relative 3D keypoints, and the global root depth. The regression training is empowered by a recent fast and strong vision transformer, Hiera [60], pre-trained with masked auto-encoder [29]. A multi-level feature fusion that concatenates the features of different layers is adopted for better feature extraction at different scales. The method additionally uses other publicly available datasets for training, namely FreiHAND [88], DexYCB [5], and CompHand [7]. The implementation is available in [83].
@@
Table 1: Method and preprocessing summary in AssemblyHands. We summarize submitted methods in terms of learning methods, architecture, preprocessing, and multi-view fusion techniques. The tuple (views, phase) indicates the number of views used in either train or test time.
@@
JHands [85] Regression Hiera [60]
@@
FRDC: This method adopts a hybrid approach by combining regression with heatmap for training. HandOccNet [53] is modified to regress 3D keypoint coordinates and integrated with an additional branch of 2D heatmap regression. HandOccNet enriches feature extraction with spatial attention mechanisms for occluded regions, making it robust under hand-object occlusions. The method further utilizes a stronger ConvNeXt [41] backbone and feature fusion from the 2D keypoint regressor.
@@
While Base uses a simple average of predicted keypoints from all four camera views in the test time, PICO-AI proposes multi-view feature fusion during training using Feature Transform Layers (FTL) [58]. This FTL training requires fusing two out of four views; thus, the method chooses the most suitable views for every frame. In cases with multiple candidates, the Intersection over Union (IoU) is computed between hand boxes from per-view predictions and 2D keypoints from previous 3D predictions. The two views with the highest IoUs are selected for their superior prediction reliability.
The methods JHands, FRDC, and Phi-AI apply adaptive fusion in predicted keypoints during testing. The method JHands computes the MPJPE with each other view and selects two results of views with the lowest MPJPE, excluding noisy predictions in the average. If the MPJPE is lower than a threshold, the mean of the two results is calculated as the final result. Otherwise, the result with a lower PA-MPJPE with the predictions in the previous frame is chosen. The methods FRDC and Phi-AI use a weighted average for each view prediction, assigning weights based on each view’s validation performance.
Postprocessing: Several postprocessing techniques, including test-time augmentation, smoothing, and model ensembling are used to enhance inference outcomes. In particular, the method JHands applies an offline smooth (Savitzky-Golay) filter on each video sequence.
@@ 4.2 ARCTICmethods
Table 2: Method and preprocessing summary in ARCTIC. We summarize baselines on ARCTIC in terms of input dimensions, image backbones, learning rate scheduling, training epochs, batch size and the cropping used for input. Method trains 50 epochs for decoder and 36 for backbone. +Learning rate is 1e-7 to 1e-4 with linear warmup for first 5% step, and 1e-4 to 1e-7 with cosine decay for rest.
@@
ArcticNet-SF [16]: Introduced in ARCTIC, it is a single-frame baseline. It first extracts an image feature vector from the input image; then, it regresses hand and object parameters with simple MLPs. The hand and object meshes can then be extracted via $\mathcal{H}(\cdot)$ and $\mathcal{O}(\cdot)$. For more details, see [16].
JointTransformer [32]: JointTransformer enhances ArcticNet-SF by integrating a transformer decoder instead of the MLP regressors for hand and object parameter estimation. The decoder employs learned queries for the angle of each joint, the shape and translation of each hand, and the translation, rotation, and articulation of the object. It alternates between self-attention between queries and cross-attention of queries to the elements of the backbone feature map, followed by linear layers that regress the final parameters. Specifically, separate linear layers are dedicated to regressing joint angles, hand shape, hand translation, object translation, object orientation, and object articulation. The best model uses a ViT-G [78] backbone with frozen DINOv2 weights [52].
@@
UVHand: Since ArcticNet-SF only leverages a global feature vector to estimate hand and object parameters, the image features lack local context. To address this, UVHand leverages Swin-L transformer [40] to extract image features. They then further leverage Deformable DETR [86] to encode the multiple-scale feature maps. The encoded feature maps are then aggregated via self- and cross-attention before regresssing hand and object parameters.
DIGIT [15]: The method was introduced to estimate strongly interacting hands in [15]. Since ArcticNet-SF is sensitive when the hands are interacting with objects, DIGIT was extended to the ARCTIC setting. Given an image, it first estimates hand part-wise segmentation masks and object masks. The mask predictions are fused with the image features to perform parameter estimation.
Implementation details: Table 2 shows that all methods use the default cropping as in ARCTIC to crop around the object, while AmbiguousHands performs three crops (around two hands and the object). DIGIT uses the HRNet-W32 backbone [67] and trains with a batch size of 64 with the same learning rate for all iterations. UVHand takes as input a $384\times 384$ image cropped around the object and encodes it with the Swin-L transformer [40] backbone. It was trained with a batch size of 48 with a learning rate of 2e-4 for the backbone and 1e-7 for other weights. Due to computational cost, they train 50 epochs for the decoder and 36 for the backbone. JointTransformer uses ViT-G [78] backbone with frozen DINOv2 [52] weights to train with a batch size of 64. It performs a linear warmup from 1e-7 to 1e-4 in the first 5% steps and uses cosine decay from 1e-4 to 1e-7 for the rest of the steps.